Solve the problem.
Suppose that $7000 is invested in an account where interest is compounded continuously at 3.2% per year. What is the doubling time?

A)4.6 yr
B)21.7 yr
C)2.2 yr
D)0.2 yr

Answer by stanbon(75887)

Solve the problem.
Suppose that $7000 is invested in an account where interest is compounded continuously at 3.2% per year. What is the doubling time?
------------------------
A(t) = P*e^(rt)
For doubling time:
2P = P*e^(rt)
e^rt = 2
-----
Your problem:
e^(0.032t) = 2
Take the natural log of both sides to get:
0.032t = ln(2)
t = [ln(2)]/0.032
t = 21.66 years
===========================
Cheers,
Stan H.
A)4.6 yr
B)21.7 yr
C)2.2 yr
D)0.2 yr